OpenStudy (anonymous):

find the amount accumulated when $300 is deposited quarterly for 10 years at 6%/year in an account containing $1500 at the start. assume end of period deposits and compounding at the same intervals as deposits.

ganeshie8 (ganeshie8):

use the annuity formula

ganeshie8 (ganeshie8):

\(\large FV = C\bullet \frac{(1+i)^{nt}-1}{i}\)

ganeshie8 (ganeshie8):

does that look familiar ?

OpenStudy (anonymous):

oh yes!

OpenStudy (anonymous):

so c is 1500 right

ganeshie8 (ganeshie8):

Nope, C = periodic cash payments = 300

ganeshie8 (ganeshie8):

forget about 1500 for now. we can deal wid it separately

OpenStudy (anonymous):

oh my bad

ganeshie8 (ganeshie8):

its there to trick us :) figure out other variables

OpenStudy (anonymous):

\[300*\frac{ 1+0.06/12^{12*10}-1 }{ 0.06/12 }\]

OpenStudy (anonymous):

@ganeshie8 ?

ganeshie8 (ganeshie8):

are we depositing 300 every month ?

OpenStudy (anonymous):

oh no quarterly equaling 4

ganeshie8 (ganeshie8):

yes, that means \(\large n = 4\) right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

which makes \(\large i = \frac{r}{n} = \frac{0.06}{4}\)

OpenStudy (anonymous):

correct so then \[300*\frac{ 0.06/4^{4*10}-1 }{ 0.06/4 }\]

ganeshie8 (ganeshie8):

Looks good !

ganeshie8 (ganeshie8):

So thats the future value for $300 periodic payments. simplify it

OpenStudy (anonymous):

8.97?

ganeshie8 (ganeshie8):

what do u mean 8.97 ?

OpenStudy (anonymous):

whoops

ganeshie8 (ganeshie8):

$300 every quarter for 10 years became 8.97 ? ha

OpenStudy (anonymous):

hahah oh god my bad

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i tried entering into TI-84 calc lol

ganeshie8 (ganeshie8):

thats a pain in d a.. .lol use wolfram always for these lengthy calculations :)

ganeshie8 (ganeshie8):

next look at the starting $1500

OpenStudy (anonymous):

if only they'd let us on exams! haha and ok

ganeshie8 (ganeshie8):

can u find its future value ?

OpenStudy (anonymous):

whats the formula again? fv=?

ganeshie8 (ganeshie8):

1500 fixed amount, 6% annual interest rate for 10 years quarterly compounding

ganeshie8 (ganeshie8):

for fixed deposit the Future Value formula is : \(FV = PV * (1+i)^{nt}\)

ganeshie8 (ganeshie8):

\(\large FV = 1500*(1 + \frac{0.06}{4})^{4*10}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

15920.45

ganeshie8 (ganeshie8):

that looks way too much. check once agian..

OpenStudy (anonymous):

damn Calc lol let me try on alph

OpenStudy (anonymous):

2721.02

ganeshie8 (ganeshie8):

Correct ! add this to the previous money. total = ?

OpenStudy (anonymous):

2721.02+48841.10=51562.12

OpenStudy (anonymous):

awesome so that'd be the answer?

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

awesome, how i hate math at times haha

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

ikr.. financial algebra can make u feel that lol...

ganeshie8 (ganeshie8):

but actually only handful of formulas are there to remember

OpenStudy (anonymous):

haha yeah formulas are everywhere

ganeshie8 (ganeshie8):

if u memorize them and are able to pull them at will when required... u wil feel comfortable

ganeshie8 (ganeshie8):

have fun :)

OpenStudy (anonymous):

thank you ill definitely be studying them

ganeshie8 (ganeshie8):

np :)